OpenStudy (anonymous):

A patient with high blood pressure who weighs 226 pounds is put on a diet to lose 26 pounds in 3 months. The patient loses 6 3/4 pounds the first month and 111/2 pounds the second month. How much weight must be lost the third month for the goal to be achieved? ?lb

OpenStudy (nirmalnema):

6(3/4)=27/4 11(1/2) =23/2 27/4+23/2 (27+46)/4 73/4 in 2 months// 26-73/4 is the amount of weight more to lose (104-73)/4 =31/4 =7(3/4)

OpenStudy (nirmalnema):

@harz360 do u understand..!!

OpenStudy (anonymous):

yeah i do understand now I was doing it the wrong way thanks so much :)

OpenStudy (nirmalnema):

ok ur welcome.. :)